OpenStudy (anonymous):

Can someone please help me with this?? i will medal AND fan!!:) A ferris wheel has a radius of 80ft. Two particular cars are located such that the central angle between them is 165°. To the nearest tenth, what is the length of the intercepted arc between those two cars on the Ferris wheel?

hero (hero):

\(s = r\theta\)

hero (hero):

\(r = 80 \\\theta = 165^{\circ}\) But convert \(\theta\) to radians first before performing the multiplication